If the product of two integers is 320 and their least common multiple is 80, what is their greatest common divisor?

a) 2

b) 4

c) 5

d) 8

e) cannot be determined
Answer #1

Let the 2 integers be a and b respectively. Since the product of two integers is 320, hence ab = 320. Further, since the least common multiple is 80, hence 80 is the smallest number which is divisible by both a and b. Also, the highest value of either a or, b is 80.

Further, 80 = 24 * 5 and 320 =26 * 5 . Since 5 is a factor of 80, one of the numbers a and b must be a multiple of 5 and the other number must be a multiple of 2. Also, 320/80 = 4.

With this analysis, it is apparent that the two numbers are 4 and 80. Their product is 320 and their least common multiple is 80.

The greatest common divisor of 4 and 80 is 4.

Option b) is the correct answer.